# Turning Trash into Treasure: Accelerating Inference of Large Language Models with Token Recycling

### v1

@@ 3.3 Verification and Update
Once the merged sequence $S$ is constructed, it is combined with the pre-processed attention mask into the LLM for a forward pass. The verification aligns with Section 2.3. All potential draft sequences are extracted based on $Tree$, and the longest correct sequence is selected as the final output.
@@
It is important to note that duplicates may be in $S$ since multiple preceding tokens may have the same candidate token. Consequently, the output distributions of duplicates are likely to differ. When performing updates in parallel, CUDA operations may merge these updates, leading to variations in the final result. For example, if $x_{i}$ appears twice and has two different top-2 output tokens, $[y_{0},y_{1}],[z_{0},z_{1}]$, then $\mathcal{M}[x_{i}]$ could be updated to exactly one of the following results: $[y_{0},z_{1}]$, $[y_{0},y_{1}]$, $[z_{0},z_{1}]$ or $[z_{0},y_{1}]$. We do not resolve this merging, as adding controls reduces overall performance, as discussed later in Section 5.2.
The update process directly overwrites the previous candidate tokens, and leverages the new ones as draft tokens for subsequent decoding steps. This allows the retrieval space to dynamically adapt to the current content, focusing on the most relevant and probable continuations. It also eliminates the necessity for extra operations beyond the standard decoding to update the retrieval space.
@@
Medusa* 1.90 1.57 1.48 1.58 1.87 1.45 2.31 89.41 1.65 - - -
Ours 2.17 1.90 1.94 1.95 2.40 1.78 2.70 110.06 2.03 2.93 131.20 2.34
shuffle 2.16 1.91 1.93 1.94 2.37 1.72 2.71 109.07 2.01 - - -
13B Vanilla 1.00 1.00 1.00 1.00 1.00 1.00 1.00 39.41 1.00 1.00 41.31 1.00
Lookahead 1.29 1.06 1.16 1.12 1.48 1.09 1.63 47.50 1.21 1.73 56.87 1.38
@@
Medusa* 1.94 1.66 1.57 1.62 1.98 1.53 2.39 67.92 1.72 - - -
Ours 1.98 1.77 1.89 1.75 2.21 1.73 2.72 74.57 1.89 3.08 93.42 2.26
33B Vanilla 1.00 1.00 1.00 1.00 1.00 1.00 1.00 18.44 1.00 1.00 19.44 1.00
Lookahead 1.32 1.09 1.20 1.17 1.55 1.14 1.61 23.03 1.25 1.70 29.22 1.50
@@
Medusa* 1.98 1.75 1.63 1.68 2.09 1.61 2.32 33.11 1.80 - - -
Ours 1.95 1.75 1.92 1.77 2.24 1.78 2.63 35.16 1.91 3.05 45.43 2.34
Table 1: Performance comparison of different methods on SpecBench (Vicuna) and on MBPP (Code Llama) across three parameter sizes. Vanilla represents AR decoding. Metrics include Mean Accepted Token (MAT), Tokens per Second (Ts/s), and Speedup ratio relative to Vanilla (Speed). Categories for SpecBench include cover Multi-turn Conversation (MT), Translation (Trans), Summarization (Sum), Question Answering (QA), Mathematical Reasoning (Math), and Retrieval-Augmented Generation (RAG), with Speed as the displayed metric. MBPP results exclude Medusa as it lacks a Code Llama variant. Medusa* involves training while others are training-free. shuffle indicates performance after shuffling the data order (Section 5.2).
In summary, Token Recycling capitalizes on the “trash” present in speculative decoding by implementing a cycling process between candidate and draft tokens. It accelerates inference without the need for additional model structures or training, making it highly adaptable and seamlessly integrated with any architecture or model size.
@@ 4.1 Experimental Setup
Align with previous work  (Kou et al. 2024), we focus on common computational redundancy scenarios, specifically greedy decoding with a batch size of one. The following evaluation metrics are used: Mean Accepted Token (MAT) (Xia et al. 2024) represents the average number of tokens confirmed in a single decoding step; Tokens per Second (Ts/s) measures the number of tokens processed per second; Speedup ratio compares the performance relative to HuggingFace’s implementation of AR decoding. We set $k=8$ for $\mathcal{M}$ (<2MB storage in sum) and the draft tree structure is shown in Appendix 8.1. All experiments are conducted using Pytorch 2.3 with a single A100-80GB GPU and 128 CPUs under CUDA 12.2.
@@
We conduct experiments on SpecBench (Xia et al. 2024) and MBPP (Austin et al. 2021). SpecBench is a comprehensive benchmark encompassing diverse scenarios including Multi-turn Conversation, Translation, Summarization, Question Answering, Mathematical Reasoning, and Retrieval-Augmented Generation. MBPP is a widely used dataset in code generation, a domain with a growing demand for efficient generation. These datasets enable a comparative analysis with prior work across both general and specialized domains. For the LLMs, we follow the standard practice of utilizing Vicuna (Chiang et al. 2023) for SpecBench and Code Llama (Roziere et al. 2023) for MBPP across three different scales: 7B, 13B, and 33B [^1] The largest model of Code Llama is 34B, for consistency and convenience in our comparisons, we refer to it as 33B..
@@
We compare Token Recycling with three lossless, train-free, and retrieval-based methods. Lookahead constructs an n-gram retrieval library through additional n-gram generation during decoding, consuming significant computational resources. PLD treats previous content as the retrieval library, which is constrained and cannot introduce new tokens or new token combinations. REST builds the retrieval library from existing training datasets, requiring large storage and considerable retrieval time. The static nature of the library also prevents it from adapting to individual queries. Furthermore, we include a parameter-efficient tuning baseline for border comparison. Medusa adds multiple additional LM heads in the final layer to predict draft tokens. We focus on losses Medusa-1 since Medusa-2 is lossy. All baselines use their default hyperparameters.
@@ 4.2 Main Results
Figure 3: Effects of tree breadth and depth on MAT and Tokens/s in (a) and (b), and the impact of updating strategies in (c).
@@
Table 1 shows the performance of Token Recycling compared to other methods. On SpecBench, it achieves more than a 2x speedup on the 7B model, nearly 30% higher than the previous train-free methods. Even compared to tuning Medusa, it shows an improvement of almost 25%. For the 13B and 33B models, it consistently provides nearly 2x speedup, maintaining the 30% acceleration advantage. These results demonstrate that Token Recycling is the most effective train-free method on SpecBench, offering substantial and consistent speedup across all model sizes.
@@
Notably, Token Recycling achieves the best speedup across most sub-tasks as well, except it slightly trails PLD on Summarization. This may be due to this task often involves many repetitions of previous content. However, the performance gap between Token Recycling and PLD narrows as the model size increases, reaching only a 1% difference with the 33B model. This is due to larger models tending to generate new tokens rather than repeat previous content. In other tasks such as Multi-turn Conversation, Translation, Question Answering, and Mathematical Reasoning, Token Recycling shows a significant improvement of about 40%~70% for the 7B model. This demonstrates the strong generalization of our method across various scenarios. Although the improvement on Retrieval-Augmented Generation is less than 3% for the 7B model, it increases with model size, exceeding 10% for the 33B one. This improvement is consistent with the preference of larger models for new tokens. Compared to the general domain, all methods achieve greater acceleration on the code domain due to its higher content redundancy. Token Recycling provides approximately 2.3x speedup across all model scales, achieving the SOTA performance.
@@
Furthermore, performances on Trans show the advantages of our method compared to PLD and REST. While PLD shows negligible speedup (close to 1x) and REST achieves its lowest speedup across tasks, Token Recycling consistently delivers over 1.75x speedup across all model sizes. Notably, on the 7B model, PLD results in a slowdown, and REST achieves just 1.05x, whereas Token Recycling reaches 1.9x. Trans requires generating new tokens continuously, involving minimal repetition of previous content. Additionally, it is highly context-sensitive, making it challenging to find exact matches from any pre-existing database. These pose challenges for PLD and REST. In contrast, the adaptive and diverse retrieval space of Token Recycling leads to superior performance. In addition to Speed, Token Recycling achieves the highest MAT across both benchmarks. This is attributed to its shorter retrieval times and the avoidance of additional generations like Lookahead. This allows for deeper and wider draft trees, enabling more tokens to be accepted in a single decoding step.
@@
Token Recycling demonstrates significant improvements across all scenarios, highlighting its efficiency and broad applicability. Importantly, our method is train-free and self-drafting, allowing for an approximate 2x speedup that can be seamlessly applied as a “free lunch” to any existing LLM.
@@ 5.1 Tree Structure
As previously outlined in Section 3.2, our tree structure is static and imbalanced. The tree size is a crucial factor to accelerate. A larger tree allows more tokens confirmed in one decoding step but also introduces more computational overhead, increasing the time required for each decoding step. To investigate the impact of tree size, specifically its depth and breadth on acceleration performance, experiments are conducted on MT-Bench using Vicuna-7B.
@@
Increasing the breadth of the tree allows for covering more possibilities during decoding. In Figure 3(a), the breadth is expanded by adding nodes while keeping the depth fixed at six layers. This leads to a consistent improvement in MAT. However, when the breadth exceeds 80, Tokens/s begins to decrease. The additional computational overhead eventually outweighs the benefits of a higher MAT.
@@ 5.2 Ablation Study
Data Order
In Token Recycling, the adjacency matrix is initialized with hot start, meaning it inherits the matrix from the previous query. By default, SpecBench organizes data so that instances from the same domain are processed consecutively. This sequential ordering might allow domain-specific patterns to persist, potentially optimizing token retrieval and boosting performance. To evaluate the robustness of hot start, we conduct an experiment where all data instances are shuffled across different domains. This shuffling ensures that each query is independent, removing any potential domain influence from the previous query. As shown in Table 1 under shuffle, the performance impact is minimal, with the Speed decreasing by less than 1%. This demonstrates that hot start is effective and robust, even without domain continuity between queries.

### v2

@@ 3.3 Verification and Update
The verification of the draft tree aligns with Section 2.2. Merged sequence $S$ is constructed through traversing the draft tree by layers. All potential draft sequences are then verified and the longest correct sequence is selected.
@@
Since multiple preceding tokens may have the same candidate token, duplicates may appear in $S$, and their output distributions are likely to differ. When performing updates in parallel, CUDA operations may merge these updates, leading to variations in the final result. For example, if $x_{i}$ appears twice and has two different top-2 output tokens, $[y_{0},y_{1}],[z_{0},z_{1}]$, then $\mathcal{M}[x_{i}]$ could be updated to exactly one of the following results: $[y_{0},z_{1}]$, $[y_{0},y_{1}]$, $[z_{0},z_{1}]$ or $[z_{0},y_{1}]$. We do not resolve this merging, as adding controls reduces overall performance, as discussed later in Section 5.2.
The update process directly overwrites the previous candidate tokens and leverages the new ones as draft tokens for subsequent decoding steps. This allows the retrieval space to dynamically adapt to the current content, focusing on the most relevant and probable continuations. It also eliminates the necessity for extra operations beyond the standard decoding to update the retrieval space.
@@
Medusa 1.90 1.57 1.48 1.58 1.87 1.45 2.31 89.41 1.65 - - -
TR 2.17 1.90 1.94 1.95 2.40 1.78 2.70 110.06 2.03 2.93 131.20 2.34
13B AR 1.00 1.00 1.00 1.00 1.00 1.00 1.00 39.41 1.00 1.00 41.31 1.00
Lade 1.29 1.06 1.16 1.12 1.48 1.09 1.63 47.50 1.21 1.73 56.87 1.38
@@
Medusa 1.94 1.66 1.57 1.62 1.98 1.53 2.39 67.92 1.72 - - -
TR 1.98 1.77 1.89 1.75 2.21 1.73 2.72 74.57 1.89 3.08 93.42 2.26
33B AR 1.00 1.00 1.00 1.00 1.00 1.00 1.00 18.44 1.00 1.00 19.44 1.00
Lade 1.32 1.09 1.20 1.17 1.55 1.14 1.61 23.03 1.25 1.70 29.22 1.50
@@
Medusa 1.98 1.75 1.63 1.68 2.09 1.61 2.32 33.11 1.80 - - -
TR 1.95 1.75 1.92 1.77 2.24 1.78 2.63 35.16 1.91 3.05 45.43 2.34
Table 1: Performance of different methods on SpecBench (Vicuna) and on MBPP (Code Llama) across all parameter sizes. Speed is the displayed metric for categories of SpecBench. MBPP results exclude Medusa as it lacks a Code Llama variant. Medusa involves training while others are training-free. Bold represents the highest performance.
In summary, TR capitalizes on the ‘trash’ present in speculative decoding by implementing a cycling process between candidate and draft tokens. It accelerates inference without the need for additional model structures or training, making it highly adaptable and seamlessly integrated with any architecture or model size.
@@ 4.1 Experimental Setup
Align with previous work  Kou et al. (2024), we focus on common computational redundancy scenarios, specifically greedy decoding with a batch size of one. The following evaluation metrics are used: Mean Accepted Token (MAT) Xia et al. (2024) represents the average number of tokens confirmed in a single decoding step; Tokens per Second (Ts/s) measures the number of tokens processed per second; Speedup ratio compares the performance relative to HuggingFace’s implementation of AR decoding. We set $k=8$ for $\mathcal{M}$ (<2MB storage in sum) and the draft tree structure is shown in Appendix 1. All experiments are conducted using Pytorch 2.3 with a single A100-80GB GPU and 128 CPUs under CUDA 12.2.
@@
We conduct experiments on SpecBench Xia et al. (2024) and MBPP Austin et al. (2021). SpecBench is a comprehensive benchmark encompassing diverse scenarios including Multi-turn Conversation (MT), Translation (Trans), Summarization (Sum), Question Answering (QA), Mathematical Reasoning (Math), and Retrieval-Augmented Generation (RAG). MBPP is a widely used dataset in code generation, which has a growing demand for efficient generation. These datasets enable a comparative analysis with prior work across both general and specialized domains. We follow the standard practice of utilizing Vicuna Chiang et al. (2023) for SpecBench and Code Llama Roziere et al. (2023) for MBPP across three different scales: 7B, 13B, and 33B [^1] The largest model of Code Llama is 34B, for consistency and convenience in our comparisons, we refer to it as 33B..
@@
We compare TR with three train-free retrieval-based methods. Lookahead (Lade) constructs an n-gram retrieval library through additional n-gram generation during decoding, consuming significant computational resources. PLD treats previous content as the retrieval library, which is constrained and cannot introduce new tokens or new token combinations. REST builds the retrieval library from existing training datasets, requiring large storage and considerable retrieval time. The static nature of the library also prevents it from adapting to individual queries. Furthermore, we also include a train-need baseline for border comparison. Medusa adds multiple additional LM heads in the final layer to predict draft tokens. We focus on losses Medusa-1 since Medusa-2 is lossy. All baselines use their default hyperparameters.
@@ 4.2 Main Results
Method Memory (MB) Speed
Lade 105 1.27
PLD 0 1.53
REST 465 1.22
Medusa >800 1.65
TR 1.95 2.03
Table 2: The additional memory costs for all methods. Medusa adds extra LM heads to the model, so the memory usage depends on the hidden size and the precision. 800MB is based on a 7B LLM and fp16 precision.
@@
Table 1 shows the performance of TR compared to other methods. On SpecBench, it achieves more than a 2x speedup on the 7B model, nearly 30% higher than the previous train-free methods. Even compared to tuning Medusa, it shows an improvement of almost 25%. For the 13B and 33B models, it consistently provides nearly 2x speedup, maintaining the 30% acceleration advantage. These results demonstrate that TR is the most effective train-free method on SpecBench, offering substantial and consistent speedup across all model sizes.
@@
Notably, TR achieves the best speedup across most sub-tasks as well, except it slightly trails PLD on Sum. This may be due to this task often involves many repetitions of previous content. However, the performance gap between TR and PLD narrows as the model size increases, reaching only a 1% difference with the 33B model. This is due to larger models tending to generate new tokens rather than repeat previous content. In other tasks such as MT, Trans, QA, and Math, TR shows a significant improvement of about 40%~70% for the 7B model. This demonstrates the strong generalization of our method across various scenarios. Although the improvement on RAG is less than 3% for the 7B model, it increases with model size, exceeding 10% for the 33B one. This improvement is consistent with the preference of larger models for new tokens. Compared to the general domain, all methods achieve greater acceleration on the code domain due to its higher content redundancy. TR provides approximately 2.3x speedup across all model scales, achieving the SOTA performance.
Figure 3: Effects of tree breadth, depth and updating strategies on MAT and Tokens/s are in (a), (b), and (c).
@@
Furthermore, performances on Trans show the advantages of our method compared to PLD and REST. While PLD shows negligible speedup (close to 1x) and REST achieves its lowest speedup across tasks, TR consistently delivers over 1.75x speedup across all model sizes. Notably, on the 7B model, PLD results in a slowdown, and REST achieves just 1.05x, whereas TR reaches 1.9x. Trans requires generating new tokens continuously, involving minimal repetition of previous content. Additionally, it is highly context-sensitive, making it challenging to find exact matches from any pre-existing database. These pose challenges for PLD and REST. In contrast, the adaptive and diverse retrieval space of TR leads to superior performance. In addition to Speed, TR achieves the highest MAT across both benchmarks. This is attributed to its shorter retrieval times and the avoidance of additional generations like Lade. This allows for deeper and wider draft trees, enabling more tokens to be accepted in a single decoding step.
Tokens/s Speed
AR 54.98 1.00
Random 95.07 1.73
Zero 102.68 1.87
Fixed 117.43 2.12
Shuffle 118.78 2.16
TR 119.56 2.17
Table 3: The impact of different initialization strategies of the adjacency matrix. Random means randomly selected from the vocabulary, Zero means all set to zero, Fixed means inherited from a fixed matrix and Shuffle means shuffle the test set.
Table 2 summarizes the GPU memory requirement for all methods. Compared to REST and Lade, TR achieves higher speedup with far less memory. While PLD requires no additional memory, its speedup is limited. Unlike Medusa, our approach is training-free, requires minimal memory, and still achieves superior performance.
@@
TR demonstrates significant improvements across all scenarios, highlighting its efficiency and broad applicability. Importantly, TR is train-free and self-drafting, allowing for an approximate 2x speedup that can be seamlessly applied as a ‘free lunch’ to any existing LLM.
@@ 5.1 Tree Structure
As previously outlined in Section 3.2, our tree structure is static and imbalanced. The tree size is a crucial factor to accelerate. A larger tree allows more tokens confirmed in one decoding step but also introduces more computational overhead, increasing the time required for each decoding step. To investigate the impact of tree size, specifically its depth and breadth, experiments are conducted on MT-Bench using Vicuna-7B.
@@
Increasing the breadth of the tree allows for covering more possibilities. In Figure 3(a), the breadth is expanded by adding nodes while keeping the depth fixed at six layers. This leads to a consistent improvement in MAT. However, when the breadth exceeds 80, Tokens/s begins to decrease. The additional computational overhead eventually outweighs the benefits of a higher MAT.
@@ 5.2 Ablation Study
Hot Start
In TR, the adjacency matrix inherits from the previous one. In Table 3, we explore the impact of different initialization strategies. Random means randomly selecting tokens from the vocabulary, while Zero sets all matrix elements to zero. Fixed selects 100 queries from AlpacaEval Li et al. (2023) (unrelated to the test set), executes them, and stores the resulting matrix. This matrix is then used to initialize each query in the test set. Shuffle refers to shuffling the test set. Compared to the Zero, the irrelevant noise introduced by Random leads to a sharp decrease in performance. Fixed, Shuffle and TR show significant improvements over Zero, suggesting that the prior matrix may capture common patterns that effectively assist subsequent queries. The relatively small difference among them indicates that these patterns are generalizable and not tied to specific tasks or content.
@@
Effect of Rejected Tokens
During the update, we refresh the candidate tokens for all draft tokens, including both accepted and rejected tokens. To further illustrate the significant effect of trash tokens, we compare two settings: updating only the candidates of accepted tokens versus of all draft tokens. As shown in Table 4, including candidates of rejected tokens significantly improves the MAT. This indicates that rejected tokens also carry valuable information necessary for subsequent decoding. We include a case study in Appendix A.6.
SpecBench MBPP
Only Accepted 1.63 1.99
All Draft 2.69 2.93
Table 4: Mean Accepted Token (MAT) for updating candidate tokens from only accepted or all draft tokens.
6 Related Work
Efficient inference is crucial for real-time applications and low-resource scenarios. Many strategies have been developed to reduce latency Zhou et al. (2024b). Among these, speculative decoding Chen et al. (2023); Leviathan et al. (2023); Miao et al. (2024); Xia et al. (2023) is a losses technique that predicts multiple possible continuations simultaneously. It reduces the number of decoding steps needed without compromising accuracy. Some speculative decoding methods rely on additional draft models to guess draft tokens. These typically involve using smaller models from the same series Zhao et al. (2024); Spector and Re ; Sun et al. (2023); Liu et al. (2024b); Yuan et al. (2024); Gong et al. (2024) or training new models with a shared vocabulary Leviathan et al. (2023); Chen et al. (2023); Zhou et al. (2024a); Li et al. (2024). It is worth noting that Zhao et al. (2024) also uses rejected tokens but does not include candidate tokens. Additionally, Kou et al. (2024); Wang et al. (2024b) propose training the original LLMs to enable non-aggressive decoding. While effective, these approaches require managing or training multiple models, which can be non-trivial and resource-intensive. Other methods focus on parameter-efficient structures. These approaches minimize the need for complete retraining but still require model-specific training and adaptation, limiting their scalability and general applicability Lin et al. (2024); Liu et al. (2024a).
Train-free methods construct retrieval libraries to obtain draft tokens Yang et al. (2023). Lookahead Fu et al. (2024) generates n-grams through multiple decodings, building a retrieval library that can hit multiple tokens in one step. However, it requires the LLM to generate n-grams while responding to queries, which reduces efficiency. PLD Saxena (2023) retrieves only from previous content, resulting in minimal overhead and significant speedup in high-redundancy tasks like summarization. However, it provides little acceleration for tasks requiring the generation of new content, like translation. REST He et al. (2023) constructs retrieval libraries using existing corpora and performs well in common scenarios. However, this approach requires large storage, time-consuming retrieval, and cannot adapt to each query.
Token Recycling is a train-free, retrieval-based method. It requires no additional generation, covers a broader range of possible continuations, and demands minimal storage with low retrieval costs. The update process ensures the retrieval space remains adaptable.
7 Conclusion
In this work, we introduce Token Recycling, a speculative decoding method for accelerating the inference of LLMs. It utilizes an adjacency matrix to store candidate tokens and retrieve a draft tree, which is then verified with tree attention. The matrix is updated with new candidate tokens generated during decoding. Token Recycling could be integrated seamlessly with existing LLMs and tasks. As a train-free approach, it achieves a speedup of nearly 2x with <2MB of storage, improving over 31% compared to previous train-free approaches.
Limitations
Our study is comprehensive, but has certain limitations that we plan to address in future research. In constructing the draft tree, we use a static tree structure. However, a dynamic tree could be employed instead. While dynamic trees introduce additional complexity, they allow for better adaptation to each decoding step, potentially improving performance by tailoring the tree structure to the specific requirements of each query.
Ethical Considerations
The data for the proposed methods is drawn solely from publicly accessible project resources on reputable websites, ensuring that no sensitive information is included. Moreover, all datasets and baseline models used in our experiments are also available to the public. We have taken care to acknowledge the original authors by properly citing their work.
Appendix A Appendix
A.1 Identical Probability Tokens
Method MT-Bench GSM8K
AR Decoding 6.17 35.2
Tree Attention 6.23 35.2
Table 5: Quality/Accuracy comparison of AR-Decoding and Tree Attention on MT-Bench and GSM8K. MT-Bench results are taken from  Cai et al. (2024). It shows that Tree Attention has minimal impact on both answer accuracy and quality.
Floating-point representation in the computer has precision errors, commonly known as ‘floating-point rounding errors’. Specifically, the precision of floating-point numbers is determined by the number of bits in the mantissa. In the IEEE 754 standard, the float32 type has a 23-bit mantissa, meaning the smallest representable difference is $2^{-23}$, approximately $1.19\times 10^{-7}$. The float16 type, with a 10-bit mantissa, can represent differences as small as $2^{-10}$, or about $9.77\times 10^{-4}$. If the difference between two token probabilities is smaller than the precision limit of floating-point representation, these two probabilities will be rounded to the same value, and these tokens will be treated as having identical probabilities during sampling.
AR Decoding uses ‘torch.argmax’ to return the token with the highest probability. When the probabilities are the same, ‘torch.argmax’ defaults to returning the one with the smallest index. In Tree Attention, the number of mask tokens is increased compared to AR Decoding, and the attention score of the mask tokens after the softmax operation is not strictly zero, but rather a very small value close to zero. These tiny non-zero values perturb the hidden representations, causing tokens that originally had identical probabilities to now differ slightly, resulting in a different argmax outcome compared to AR Decoding.
Nevertheless, as shown in Table 5, due to the extremely rare occurrence of this issue and the affected probabilities being so close to each other, the impact on experimental accuracy and model performance is negligible.
A.2 Draft Tree Algorithm and Structure
Algorithm 1 Static Tree Based BFS
0:  Adjacency matrix $\mathcal{M}$, Static tree structure $Tree$, the last prompt token $x_{t}$
0:  Merged Sequence $S$
1:  Initialize $S\leftarrow\emptyset$
2:  Initialize $root\leftarrow x_{t}$
3:  Initialize the current layer $L\leftarrow(root)$
4:  Initialize the current depth $d\leftarrow 0$
5:  while $d<Tree.depth$ do
6:     Initialize next layer $L_{\text{next}}\leftarrow\emptyset$
7:     Get all candidate tokens of $L$ from $\mathcal{M}$ in parallel
8:     $xs=M[L]$
9:     Extract next layer tokens from $xs$ with $Tree$
10:     $L_{\text{next}}=xs[Tree[d].index]$
11:     Concatenate $S$ and $L$
12:     $S\leftarrow(S;L)$
13:     $L\leftarrow L_{\text{next}}$
14:  end while
15:  return  $S$
Utilizing tree attention  Miao et al. (2024) to extend the path in the verification phase has become a widely adopted strategy for speculative decoding methods.
In Token Recycling, we also use a heuristically constructed token tree to perform the verification. As shown in Figure 5, we construct a static and unbalanced tree inspired by  Cai et al. (2024). The number $k$ on a node indicates that it is the $k$-th candidate token for its parent node. The construction process is below. We begin with a fully balanced 10-branch tree and use an independent validation set to identify the top $k$ nodes that most frequently yield correct tokens. These top $k$ nodes and their children are retained to form a new tree, and the process is repeated to identify the next set of top $k$ nodes. This iterative process continues until performance no longer shows significant improvement. The final tree is determined, and the $k$ is set to consider the maximum number of children across all nodes and the memory requirement. While empirical, this iterative approach has proven to be effective. Further details on tuning the $n$ are provided in Section 5.1. Overall, the tree we construct contains 80 nodes (including the root node) in 6 layers. This means that each forward requires an additional draft input of 79 tokens with a maximum acceptance length of 6.
Building on the tree structure described above, we construct a draft tree for the current content by a BFS-like algorithm in the inference phase. As described in Algorithm 1, we infill the child nodes of each layer in turn according to the matrix. At last, the merged sequence $S$ is returned and sent to tree attention with $Tree$.
A.3 Time Allocation